OpenStudy (anonymous):

can anyone help me i will meadel and fan solve the equations 1. 3x-5=13 2. R/2 - 6=14 3. 4m + 12 = 24

OpenStudy (adi3):

all 3

OpenStudy (adi3):

together or like the 1st equation first then the second

OpenStudy (anonymous):

no it is not right because the answer chocies dosent say that man the answer chocies are diffrent

OpenStudy (adi3):

read my question should we solve all 3 together or together or like the 1st equation first then the second

OpenStudy (anonymous):

well for the first, you have to first add 5 to 13 so 3x=18

OpenStudy (anonymous):

oh first second and third

OpenStudy (anonymous):

sorry about that adi3

OpenStudy (adi3):

no prob

OpenStudy (anonymous):

r/2=20 r=40 4m =12 m =3

OpenStudy (adi3):

@tylermcmullen23 is right for 1st add 5 + 13 = 18

OpenStudy (adi3):

5 = 18 5/18 =

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

that should help with the other 2, you just follow the same steps. the number without the variable is the one you add/subtract to the number on the other side of the "="

OpenStudy (adi3):

got it then medal someone